For the reaction below the Kc is 2.24 x10^22. What is the Kc for the reverse reaction? 2Co(g)+ O2(g)=2 CO2(g)
Answer #1

Solution :-

Kc for the forward reaction is 2.24*10^22

Kc for the reaverse reaction = ?

when the reaction is reversed then the Kc value is inversed

so

Kc reverse = 1/ Kc forward

                 = 1/ 2.24*10^22

                 = 4.46*10^-23

So the Kc for the reaverse reaction is 4.46*10^-23